### v1

@@ 4.2 Main convergence results
Let $x\in\Omega$ be a feasible point for the MPEC (1), which is not B-stationary in the sense of Definition 2.1. Then there exists an $\epsilon>0$, a step $s\in{\mathbb{R}}^{n}$ with $\|s\|_{\infty}=1$ and a partition $({\mathcal{D}}_{1}(x),{\mathcal{D}}_{2}(x))$ of ${\mathcal{I}}_{00}(x)$ such that:
@@
$\nabla c_{i}(x)^{\top}s\geq\epsilon$ $i\in{\mathcal{A}}(x),$ (28b)
@@
Proof. If $x$ is not a B-stationary point then it follows from (10) that it is not stationary for at least one BNLP. The assertion follows by regarding every corresponding LP piece. Since $d=0$ is not a solution for these LPs, then there exists a descent direction $s$ on at least one BNLP. ∎
Next, we show that an LPEC constructed at an almost feasible point can predict a feasible BNLP. In our implementation, such a point, and the corresponding LPEC are found via Algorithm 2.
@@
Let Assumptions 4.1-4.4 hold. Consider $x^{*}\in\Omega$ a feasible point of the MPEC (1), and $\tilde{x}\in X$ (not necessarily an element of $\Omega$) such that $\|\tilde{x}-x^{*}\|\leq\delta$. Then, there exist positive constants $\kappa$, $\epsilon$ such that for a sufficiently small $\delta$ the following interval is nonempty
@@
Proof. Consider the case when $\tilde{x}\in\Omega$, i.e., $\delta=0$. The $\mathrm{LPEC}(\tilde{x},\rho)$ is always feasible for this point and let $d$ be an element of its feasible set. Now set $\kappa=\bar{\rho}$, where $\bar{\rho}$ is defined in (14). Then according to Lemma 3.2, for all $\rho\in(0,\kappa)$, every ${\mathrm{BNLP}}({\mathcal{I}}_{1}(\tilde{x}+d),{\mathcal{I}}_{2}(\tilde{x}+d))$ is feasible.
@@
$\nabla f(\tilde{x})^{\top}d<-\epsilon,\nabla c_{i}(\tilde{x})^{\top}s\geq 0,\ i\in{\mathcal{A}}(x^{*}).$ (30)
Next, we prove that every ${\mathrm{LPEC}}(\tilde{x},\rho)$ for all $\tilde{x}\in\mathcal{N}^{*}$ and a suitable $\rho$ is feasible. We start first with the general inequality constraints and use an argument similar to (Chin2003, , Lemma 4). Take the step $d=\rho s$, which satisfies $\|d\|\leq\rho$, and consider the lineraization of the inequality constraints active at $x^{*}$:
$c(\tilde{x})+\rho\nabla c_{i}(\tilde{x})^{\top}s\geq-h_{c}(\tilde{x})+\rho\epsilon,$ $i\in{\mathcal{A}}(x^{*}).$ (31)
The first term is lower bounded by the infeasibility of the inequality constraints, the second uses the fact that $\|s\|_{\infty}=1$ and (30). For the inequality constraints inactive at $x^{*}$, it follow from the continuity of $c$ that there exists positive constants $\gamma_{1}$ and $\gamma_{2}$ independent of $\tilde{x}$ such that $c(\tilde{x})\geq\gamma_{1}$ and $\nabla c_{i}(\tilde{x})^{\top}s\geq-\gamma_{2}$. Thus for the inactive constraints, we have
$c(\tilde{x})+\rho\nabla c_{i}(\tilde{x})^{\top}s\geq\gamma_{1}-\rho\gamma_{2},$ $i\notin{\mathcal{A}}(x^{*}).$ (32)
@@
The lower bound is obtained if the complementarity constraint is satisfied by always settings the smaller component of $(\tilde{x}_{1,i},\tilde{x}_{2,i})$ to zero via the step $d$. Similarly, the upper bound is obtained by setting the larger component of $(\tilde{x}_{1,i},\tilde{x}_{2,i})$ to zero via the step $d$. The lower bound is equal to the complementarity residual $h_{\perp}(\tilde{x})$. Moreover, from Lipschitz continuity if follows that $h_{\perp}(\tilde{x})\leq\delta$ and $h_{c}(\tilde{x})\leq\delta$. Because, $\tilde{x}\in X$, there exists a constant $\kappa_{2}>0$ independent of $\tilde{x}$ such that $\|\tilde{x}\|\leq\kappa_{2}$. Therefore, for the feasibility of the complementarity constraints we require that:
@@
Furthermore, we conclude the following about the feasible ${\mathrm{LPEC}}(\tilde{x},\rho)$. Let $\tilde{d}$ be any feasible point of ${\mathrm{LPEC}}(\tilde{x},\rho)$. Due to (36), all points $\tilde{x}+\tilde{d}$ that can be reached from $\tilde{x}$ and that satisfy the complementarity constraints, can also be reached via $x^{*}+d$, where $d$ is a feasible point of ${\mathrm{LPEC}}(x^{*},\bar{\rho})$. Because Lemma 3.2 holds for ${\mathrm{LPEC}}(x^{*},\bar{\rho})$, we conclude that also all ${\mathrm{BNLP}}({\mathcal{I}}_{1}(\tilde{x}+\tilde{d}),{\mathcal{I}}_{2}(\tilde{x}+\tilde{d}))$ are feasible. This argument is illustrated in Figure 5. This concludes the proof. ∎
@@
We apply this result now to Algorithm 2.
@@
Let Assumptions 4.1-4.4 hold. Let $x^{*}(\tau)$ be a stationary point of $\mathrm{Reg}(\tau)$ for some $\tau>0$. There exist positive constants $\bar{\tau}$, $\kappa$ such that for any: $\tau\in\left(0,\bar{\tau}\right]$ and $\rho\in[\sqrt{\tau},\kappa]$ the $\mathrm{LPEC}(x^{*}(\tau),\rho)$ in (13) is feasible. Moreover, for any feasible point $d$ of this LPEC, it follows that $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau)+d),{\mathcal{I}}_{2}(x^{*}(\tau)+d))$ is feasible as well.
@@
Second, for $x^{*}(\tau)\notin\Omega$ set $\tilde{x}=x^{*}(\bar{\tau})$ and apply Theorem 4.1. Since $x^{*}(\bar{\tau})$ is a stationary point of $\mathrm{Reg}(\bar{\tau})$, it holds that $c(x^{*}(\bar{\tau}))\geq 0$, i.e., $h_{c}(\tilde{x})=0$. Moreover, form $x_{1,i}x_{2,i}\leq\tau,\ \forall i\in\{1,\ldots,m\}$, it follows that $h(\tilde{x})\leq\sqrt{\bar{\tau}}=\delta$. Hence, for a sufficiently small $\delta=\bar{\tau}$ and every $0<\tau\leq\bar{\tau}$ the interval in (37) is nonempty and we can apply Theorem 4.1.
Remarkably, the LPEC in the regularization-based Phase I does not even need to be solved for local optimality. It is sufficient to find a feasible point of the LPEC to identify a feasible BNLP. Now we have all the ingredients to state our main convergence result.
@@
Regard the MPEC (1) and suppose that Assumptions 4.1-4.4 hold. Then MPECopt given in Algorithm 3, with a sufficiently large finite number of iterations $N^{\mathrm{out}}$ and $N^{\mathrm{in}}$, with a Phase I from Section 3.3 has one of the following mutually exclusive outcomes:
@@
    It concludes that the MPEC (1) is locally infeasible.
@@
Proof. We discuss Phase I first. Consider Algorithm 2. It follows from Lemma 3.5 that if $\mathrm{Reg}(\tau)$ is locally infeasible, then so is the MPEC  (1). On the other hand, if the MPEC is feasible, then it follows from Theorem 4.1 that Algorithm 2 will eventually find a feasible point $x^{0}$. Similarly, it follows from the discussion in Section 3.3.2, by solving the $\ell_{1}$ or $\ell_{\infty}$ reformulation of (25), local infeasibility is certified or a feasible point $x^{0}$ is found. This completes the proof for the first outcome.
Next, given a feasible point $x^{0}$, we show the finite termination of the Algorithm 3. Consider some iterate $x^{k}$, which is a stationary point of some $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}),{\mathcal{I}}_{2}(x^{k}))$, and a $\rho^{k,l}>0$. If $d^{k,l}=0$ is a local minimizer of $\mathrm{LPEC}(x^{k},\rho^{k,l})$, then a B-stationary point is found and the algorithm terminates. Similarly, if ${\mathcal{I}}_{00}(x^{k})$ is empty, then $x^{k}$ is B-stationary. Otherwise, the $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$ is solved for $x^{k,l}$. If $f(x^{k,l})<f(x^{k})$ the step is accepted and $x^{k+1}=x^{k,l}$. If not, the radius of the trust region is reduced.
Now by Lemma 3.1 for a sufficiently small $\rho^{k,l}$ the full (13) and reduced LPECs (12) have the same set of local minimizers. Next, it follows from Lemma 3.2 that the point $x^{k}$ is feasible for any $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$. It follows from the equation (1) that a local minimizer of an LPEC is also a local minimizer for all branch LPs. Consider the $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$ and its corresponding branch LP, whose solution is $d^{k,l}\neq 0$. This means that $x^{k}$, which is feasible for $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$, is not a stationary point of this NLP, and according to the Assumption 4.1, the NLP solver will return a point with a lower objective value. From this analysis, we conclude that if $d^{k,l}\neq 0$ is a local minimizer of $\mathrm{LPEC}(x^{k},\rho^{k,l})$, then an acceptable step will be computed for a sufficiently small $\rho^{k,l}>0$.
This procedure generates a sequence of $\Delta f^{k}=f^{k}-f^{k+1}>0$. However, this sequence is also finite because the number of possible BNLPs is finite, and $\Omega$ is compact. Consider the last point $x^{K}$ in this sequence, found in a finite number of steps. Because no $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{K}),{\mathcal{I}}_{2}(x^{K}))$ can produce a better objective, we conclude from Eq. (10) that $x^{K}$ stationary for all these BNLP, hence it is B-stationary for the MPEC (1). This concludes the proof. ∎

### v2

@@ 4.2 Main convergence results
Let $x\in\Omega$ be a feasible point for the MPEC (1), satisfying Assumption 4.2, and assume that $x$ is not B-stationary in the sense of Definition 2.1. Then there exist $\epsilon>0$, a step $s\in{\mathbb{R}}^{n}$ with $\|s\|_{\infty}=1$, and a partition $({\mathcal{D}}_{1}(x),{\mathcal{D}}_{2}(x))$ of ${\mathcal{I}}_{00}(x)$ such that:
@@
$\nabla c_{i}(x)^{\top}s\geq\epsilon,$ $i\in{\mathcal{A}}(x),$ (28b)
@@
Proof. If $x$ is not B-stationary, then there exists at least one BNLP for which $x$ is not stationary. Denote by $({\mathcal{D}}_{1}(x),{\mathcal{D}}_{2}(x))$ the partition of ${\mathcal{I}}_{00}(x)$ associated with this BNLP.
By Proposition 4.1, the classical MFCQ holds for this BNLP. From this we draw two conclusions.
First, since $x$ is not B-stationary for this BNLP, Theorem 2.1 implies the existence of a direction $d\in\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ such that $\nabla f(x)^{\top}d<0$. Moreover, because MFCQ holds, the tangent cone $\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ coincides with the classical linearized feasible cone of the BNLP, rendering (4) into an LP.
Second, considering the active sets for the BNLP at $x$, MFCQ implies the existence of a nonzero direction $s$ that points into the strict interior of the LP’s feasible set. In other words, the following set is nonempty:
$S=\{\,$ $s\in{\mathbb{R}}^{n}\mid\nabla c_{i}(x)^{\top}s>0,\ i\in{\mathcal{A}}(x);$
$s_{1,i}=0,\ i\in{\mathcal{I}}_{0+}(x);\ s_{2,i}=0,\ i\in{\mathcal{I}}_{+0}(x);$
$s_{1,i}=0,\ s_{2,i}>0,\ i\in{\mathcal{D}}_{1}(x);\ s_{1,i}>0,\ s_{2,i}=0,\ i\in{\mathcal{D}}_{2}(x)\,\}.$
Combining these two facts yields the existence of $\epsilon>0$ and $s$ with $\|s\|_{\infty}=1$ such that (28) holds. ∎
Next, we show that an LPEC constructed at an almost feasible point can predict a feasible BNLP. In our implementation, such a point, and the corresponding LPEC are found in Algorithm 2.
@@
Let Assumptions 4.1-4.3 hold. Consider $x^{*}\in\Omega$ a feasible point of the MPEC (1), and $\tilde{x}\in X$ (not necessarily an element of $\Omega$) such that $\|\tilde{x}-x^{*}\|\leq\delta$. Then, there exist positive constants $\kappa$, $\epsilon$ such that for a sufficiently small $\delta$ the following interval is nonempty
@@
Proof. Consider the case when $\tilde{x}\in\Omega$, i.e., $\delta=0$ and $h_{c}(\tilde{x})=h_{\perp}(\tilde{x})=0$. The $\mathrm{LPEC}(\tilde{x},\rho)$ is always feasible for this point and let $d$ be an element of its feasible set. Now set $\kappa=\bar{\rho}$, where $\bar{\rho}$ is defined in (16). Then according to Lemma 3.2, for all $\rho\in(0,\kappa)$, every ${\mathrm{BNLP}}({\mathcal{I}}_{1}(\tilde{x}+d),{\mathcal{I}}_{2}(\tilde{x}+d))$ is feasible.
@@
$\nabla f(\tilde{x})^{\top}d<-\epsilon,\nabla c_{i}(\tilde{x})^{\top}s\geq\epsilon,\ i\in{\mathcal{A}}(x^{*}).$ (30)
Next, we prove that every ${\mathrm{LPEC}}(\tilde{x},\rho)$ for all $\tilde{x}\in\mathcal{N}^{*}$ and a suitable $\rho$ is feasible. We start first with the general inequality constraints and use an argument similar to (Chin2003, , Lemma 4). Take the step $d=\rho s$, which satisfies $\|d\|\leq\rho$, and consider the linearization of the inequality constraints active at $x^{*}$:
$c_{i}(\tilde{x})+\rho\nabla c_{i}(\tilde{x})^{\top}s\geq-h_{c}(\tilde{x})+\rho\epsilon,$ $i\in{\mathcal{A}}(x^{*}).$ (31)
The first term is lower bounded by the infeasibility of the inequality constraints, the second uses the fact that $\|s\|_{\infty}=1$ and (30). For the inequality constraints inactive at $x^{*}$, it follows from the continuity of $c$ that there exists positive constants $\gamma_{1}$ and $\gamma_{2}$ independent of $\tilde{x}$ such that $c_{i}(\tilde{x})\geq\gamma_{1}$ and $\nabla c_{i}(\tilde{x})^{\top}s\geq-\gamma_{2}$. Thus for the inactive constraints, we have
$c_{i}(\tilde{x})+\rho\nabla c_{i}(\tilde{x})^{\top}s\geq\gamma_{1}-\rho\gamma_{2},$ $i\notin{\mathcal{A}}(x^{*}).$ (32)
@@
[…]
@@
Furthermore, we conclude the following about the feasible ${\mathrm{LPEC}}(\tilde{x},\rho)$. Let $\tilde{d}$ be any feasible point of ${\mathrm{LPEC}}(\tilde{x},\rho)$. Due to (36), all points $\tilde{x}+\tilde{d}$ that can be reached from $\tilde{x}$ and that satisfy the complementarity constraints, can also be reached via $x^{*}+d$, where $d$ is a feasible point of ${\mathrm{LPEC}}(x^{*},\bar{\rho})$. Because Lemma 3.2 holds for ${\mathrm{LPEC}}(x^{*},\bar{\rho})$, we conclude that also all $){\mathrm{BNLP}}({\mathcal{I}}_{1}(\tilde{x}+\tilde{d}),{\mathcal{I}}_{2}(\tilde{x}+\tilde{d}$)) are feasible. Fig. 4 illustrates the arguments. This concludes the proof. ∎
@@
Remark 4.1
Theorem 4.1 ensures that any algorithm that produces a point sufficiently close to the MPEC feasible set can be used in Phase I. Subsequently, finding a feasible point of the LPEC constructed at this point helps identify a feasible BNLP. In Section 3.3, we introduced a specific Phase I algorithm and, in the following corollary, we specialize Theorem 4.1 to this algorithm.
@@
Let Assumptions 4.1-4.3 hold. Let $x^{*}(\tau)$ be a stationary point of $\mathrm{Reg}(\tau)$ for some $\tau>0$. There exist positive constants $\bar{\tau}$, $\kappa$ such that for any: $\tau\in\left(0,\bar{\tau}\right]$ and $\rho\in[\sqrt{\tau},\kappa]$ the $\mathrm{LPEC}(x^{*}(\tau),\rho)$ in (15) is feasible. Moreover, for any feasible point $d$ of this LPEC, it follows that $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau)+d),{\mathcal{I}}_{2}(x^{*}(\tau)+d))$ is feasible as well.
@@
Second, for $x^{*}(\tau)\notin\Omega$ set $\tilde{x}=x^{*}(\bar{\tau})$ and apply Theorem 4.1. Since $x^{*}(\bar{\tau})$ is a stationary point of $\mathrm{Reg}(\bar{\tau})$, it holds that $c(x^{*}(\bar{\tau}))\geq 0$, i.e., $h_{c}(\tilde{x})=0$. Moreover, from $x_{1,i}x_{2,i}\leq\tau,\ \forall i\in\{1,\ldots,m\}$, it follows that $h(\tilde{x})\leq\sqrt{\bar{\tau}}=\delta$. Hence, for a sufficiently small $\delta=\bar{\tau}$ and every $0<\tau\leq\bar{\tau}$ the interval in (37) is nonempty and we can apply Theorem 4.1. ∎
Remarkably, it follows from the last two results, that the LPEC in Phase I does not even need to be solved to optimality. It is sufficient to find a feasible point of the LPEC to identify a feasible BNLP.
Before we present the main convergence result, we prove that the inner loop in Algorithm 3 terminates in a finite number of iterations.
Proposition 4.3
[…]
Proof. If $d^{k,l}=0$ finite termination is trivial, so consider $d^{k,l}\neq 0$. Each step rejection in lines 10 or 12 multiplies $\rho$ by $\gamma^{\mathrm{L}}\in(0,1)$, hence after finitely many iterations there is an index $N\geq l$ with $\rho^{k,N}=(\gamma^{\mathrm{L}})^{N-l}\rho^{k,l}<\bar{\rho}$. For $\rho^{k,N}<\bar{\rho}$, due to Lemma 3.1 we can focus on the reduced LPEC (12). In this case, from the proof of Lemma 3.2 we have
$\mathcal{P}(x^{k})=\mathcal{P}(x^{k}+d^{k,N}),$
so every BNLP that can be selected at $x^{k}+d^{k,N}$ can also be selected at $x^{k}$, hence $x^{k}$ is feasible for all these BNLPs.
Since no termination occurred before index $N$, it holds that $\nabla f(x^{k})^{\top}d^{k,N}<0$ and $d^{k,N}\neq 0$. By (10) the vector $d^{k,N}$ solves all branch LPs of the reduced LPEC, and by (40) each such branch LP corresponds to an LP of the form (4) for the BNLPs. Therefore $d^{k,N}\neq 0$ is a common minimizer of all these LPs with $\nabla f(x^{k})^{\top}d^{k,N}<0$, which implies that $x^{k}$ is feasible but not stationary for any BNLP indexed by $({\mathcal{I}}_{1}(x^{k}+d^{k,N}),{\mathcal{I}}_{2}(x^{k}+d^{k,N}))\in\mathcal{P}(x^{k}+d^{k,N})$.
By Assumption 4.3, the NLP solver applied to any such BNLP produces a point $x^{k,N}$ with $f(x^{k,N})<f(x^{k})$. Hence an acceptance occurs and the inner loop terminates. Since $N$ is finite, the inner loop terminates after finitely many iterations. ∎An immediate consequence of this proposition is that we can terminate an LPEC solver early at a nonzero feasible point $d\neq 0$.
Corollary 4.2
[…]
Proof. It follows from the proof of the previous proposition that, for sufficiently large $N$, the solution of the LPEC, $d^{k,N}\neq 0$, also solves every branch LP. By item (3) of Proposition 4.1, and for $\rho^{k,N}<\bar{\rho}$, the feasible set of a branch LP is a convex cone intersected with the trust-region constraints. Because it is a cone, from the feasibility of $\hat{d}=0$ and $d^{k,N}$ it follows that any $d=td^{k,N}$ with $t>0$ is also feasible (within the trust region), and $\nabla f(x^{k})^{\top}(td^{k,N})=\nabla f(x^{k})^{\top}d<0$. Moreover, because $d=t\ d^{k,N}$ it holds that $(\mathcal{I}_{1}(x^{k}+d),\mathcal{I}_{2}(x^{k}+d))$ coincides with $(\mathcal{I}_{1}(x^{k}+d^{k,N}),\mathcal{I}_{2}(x^{k}+d^{k,N}))$. ∎
@@
Regard the MPEC (1) and suppose that Assumptions 4.1-4.3 hold. Then MPECopt given in Algorithm 3, with a Phase I from Section 3.3 has one of the following mutually exclusive outcomes:
@@
    In Phase I, a stationary point of the problem of minimizing a measure of constraint infeasibility is returned.
@@
  3. 3.
    Create an infinite sequence of iterates $\{x^{k}\}$ with decreasing objective values, whose accumulation point is B-stationary.
If, in addition, every stationary point $\hat{x}$ of each BNLP is isolated, i.e., there exists a neighborhood $\hat{\mathcal{N}}$ of $\hat{x}$ containing no other stationary points, then outcome 3 cannot occur.
[…]
Given a feasible point $x^{0}$, we show the outcomes of Algorithm 3.
If follows from Proposition 4.3 that the inner loop terminates in a finite number of iterations. Therefore, the outer loop it produces a sequence of points $\{x^{k}\}\subset\Omega$ with decreasing objective values $\{f^{k}\}$.
Consider some iterate $x^{k}$, which is a stationary point of some $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}),{\mathcal{I}}_{2}(x^{k}))$, and a $\rho^{k,l}>0$. If $d^{k,l}=0$ is a minimizer of $\mathrm{LPEC}(x^{k},\rho^{k,l})$, then a B-stationary point is found and the algorithm terminates in a finite number of iterations, which is outcome two.
We prove the occurrence of outcome three. Since $\{x^{k}\}\subset\Omega\subset X$ and $f$ is continuous while $X$ is compact, the sequence $\{f(x^{k})\}$ is bounded. By monotonicity, $\{f(x^{k})\}$ is convergent to some $f^{\infty}\in{\mathbb{R}}$. Consequently, $\{x^{k}\}$ has an accumulation point $x^{\infty}\in X$. Let $\{x^{k_{j}}\}$ be a subsequence converging to $x^{\infty}$. Suppose, for contradiction, that $x^{\infty}$ is not a B-stationary point. Then, by Assumption 4.3, there exists $\alpha^{k}>0$ such that $f(x^{k})-f(x^{k+1})\;\geq\;\alpha^{k}.$ Define $\bar{\alpha}:=\min_{k}\alpha^{k}>0$. It follows that for the subsequence, $f(x^{k_{j}})-f(x^{k_{j+1}})\;\geq\;\bar{\alpha},\quad\forall j\in\mathbb{N}.$ Summing over $j$ yields the telescoping bound
$f(x^{k_{j}})=f(x^{k_{0}})-\sum_{\ell=0}^{j-1}\big(f(x^{k_{\ell}})-f(x^{k_{\ell+1}})\big)\;\leq\;f(x^{k_{0}})-j\bar{\alpha}.$
Letting $j\to\infty$ gives $f(x^{k_{j}})\to$-$\infty$. However, since $x^{k_{j}}\to x^{\infty}\in X$ and $f$ is continuous on the compact set $X$, the values $\{f(x^{k_{j}})\}$ must remain bounded. This contradiction our assumption, and therefore $x^{\infty}$ is a B-stationary point.
[…]
We note that the assumption of isolated stationary points for finite termination is not overly restrictive. Under MPEC-MFCQ in Ass. 4.2, the classical MFCQ holds for the associated BNLPs. Moreover, if we additionally assume generalized second-order sufficient conditions at the BNLP stationary points, these points are isolated local minimizers (Forsgren2002, , Theorem 6.9).